OpenStudy (anonymous):

Factor completely: 2x2 + 12x + 16 A. 2(x + 4)(x + 2) B. 2(x + 8)(x + 1) C. (x + 2)(x + 8) D. 2(x + 2)(x + 8)

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

\(2x^2 + 12x + 16 = 2 \ (x^2 + 6x + 8)\) HINT 2*4 = 8 2 + 4 = 6

OpenStudy (anonymous):

C?

OpenStudy (irishboy123):

Mmm FOIL 2(x + 4)(x + 2) = 2 ( x*x + x*2 + 4*x + 4 * 2)

OpenStudy (anonymous):

A then

OpenStudy (irishboy123):

yes!

OpenStudy (anonymous):

thanks. :)